OpenStudy (anonymous):

7(T-R) = 2R - 3(4R+T) Solve for T Can you show me steps on how to do it, or make it easier for me to understand. Thanks :)

jhonyy9 (jhonyy9):

7(T-R)= can you make the calcules here firstly ?

jhonyy9 (jhonyy9):

so will be 7(T-R)=7T-7R so the same can you make the calcule on the right side too ?

OpenStudy (anonymous):

The thing is, I can't do any of those.

jhonyy9 (jhonyy9):

so than will be 7T -7R =2R -12R -3T 7T -7R = -10R -3T 7T +3T -7R +7R = -10R +7R-3T +3T 10T = -3R T = -3R/10 ok ?

OpenStudy (anonymous):

oh, thank you. I got it now ^__^